40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 50 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  15118   Mon Jan 13 16:05:18 2020 yehonathanUpdatePSLAssembly underway for c1psl upgrade

{Yehonathan, Jon}

I configured the Acromag channels according to the Slow Controls Wiki page.

We started testing the channels. Almost at the beginning we notice that the BIO channels are inverted. High voltage when 0. 0 Voltage when 1. We checked several things:

1. We checked the configuration of the BIOs in the windows machine but nothing pointed to the problem.

2. We isolated one of the BIOs from the DIN rail but the behavior persisted.

3. We checked that the voltages that go into the Acromags are correct.

The next step is to power up an isolated Acromag directly from the power supply. This will tell us if the problem is in the chassis or the EPICs DB.

  15117   Mon Jan 13 15:47:37 2020 shrutiConfigurationComputer Scripts / Programsc1psl burt restore

[Yehonathan, Jon, Shruti]

Since the PMC would not lock, we initially burt-restored the c1psl machine to the last available shapshot (Dec 10th 2019), but it still would not lock.

Then, it was burt-restored to midnight of Dec 1st, 2019, after which it could be locked.

  15116   Fri Jan 10 19:48:46 2020 yehonathanUpdatePSLAssembly underway for c1psl upgrade

{Yehonathan, Jon}

I finished pre-wiring the PSL chassis. I mounted the Acromags on the DIN rails and labeled them. I checked that they are powered up with the right voltage +24V and that the LEDs behave as expected.

Attachment 1: 20200110_194429.jpg
20200110_194429.jpg
Attachment 2: 20200110_194516_HDR.jpg
20200110_194516_HDR.jpg
  15115   Fri Jan 10 14:21:19 2020 YehonathanUpdatePSLc1psl reboot

PSL controls on the sitemap went blank. Rebooted c1psl. PSL screens seem normal again.

  15114   Tue Jan 7 18:51:51 2020 JonUpdatePSLNew c1psl server assembled

I've assembled a new SuperMicro rackmount machine to replace c1psl. It is currently set up on the electronics bench.

  • OS: Debian 10.2
  • Hostname: c1psl1 (will become c1psl after installation)
  • IP: 192.168.113.54 (registered in the martian DNS)
  • Network drive mount point set up (/cvs/cds), which provides all the EPICS executables.
  15113   Mon Jan 6 19:05:09 2020 not gautamUpdatePSLAssembly underway for c1psl upgrade

I found them, thanks. After c1psl, there are 4 2GB DIMM cards and 1 SSD left. I moved them into the storage bins with all the other Acromag parts.

Quote:

RTFE. Where did the spares go?

Quote:

I began setting up the host server, but immediately hit a problem: We seem to have no more memory cards or solid-state drives, despite having two more SuperMicro servers. I ordered enough RAM cards and drives to finish both machines. They will hopefully arrive tomorrow.

  15112   Mon Jan 6 16:07:12 2020 gautamUpdatePSLAssembly underway for c1psl upgrade

RTFE. Where did the spares go?

Quote:

I began setting up the host server, but immediately hit a problem: We seem to have no more memory cards or solid-state drives, despite having two more SuperMicro servers. I ordered enough RAM cards and drives to finish both machines. They will hopefully arrive tomorrow.

  15111   Mon Jan 6 15:36:55 2020 JonUpdatePSLAssembly underway for c1psl upgrade

[Jon, Yehonathan]

We've begun assembling the new c1psl Acromag chassis based on Yehonathan's final pin assignments. So far, parts have been gathered and the chassis itself has been assembled.

Yehonathan is currently wiring up the chassis power and Ethernet feedthroughs, following my wiring diagram from previous assemblies. Once the Acromag units are powered, I will help configure them, assign IPs, etc. We will then turn the wiring over to Chub to complete the Acromag to breakout board wiring.

I began setting up the host server, but immediately hit a problem: We seem to have no more memory cards or solid-state drives, despite having two more SuperMicro servers. I ordered enough RAM cards and drives to finish both machines. They will hopefully arrive tomorrow.

  15110   Wed Jan 1 16:04:37 2020 YehonathanUpdatePSLMapping the PSL electronics

Done.

Quote:

For the IMC servo board, it'd be easiest to copy the wiring scheme for the BIO bits as is configured for the CM board (i.e. copy the grouping of the BIO bits on the individual Acromag units). This will enable us to use the latch code with minimal modifications (it was a pain to debug this the first time around). I don't see any major constraint in the wiring assignment that'd make this difficult.

Quote:

PSL wiring spreadsheet is ready. (But the link was stripped. Koji)

Link to a wiki page  with the link to the wiring spreadsheet (Yehonathan)

 

  15109   Wed Jan 1 14:14:00 2020 YehonathanUpdatePSLPMC Linewidth measurement

Turns out the 35MHz sidebands are way too weak to resolve from the resonance when doing a PZT scan.

I connect the IFR2023B function generator on the PSL table to the EOM instead of the FSS box and set it to generate 150MHz at 13dbm.

To observe the resulting weak sideband I place a PDA55 at the peak-off path from the transmission of the PMC where there is much more light than the transmission of the PMC head mirror. Whoever is using this path there is a PD blocking it right now.

I do a PZT scan by connecting a triangular signal to the EXT DC on the PMC servo with and without the EOM (Attachment 1). A weak sideband can clearly be spotted now.

Using the above 150MHz sideband calibration I can find the roundtrip time to be 1.55ns.

I take a high-resolution scan of a resonance peak and fit it to a Lorentzian (Attachment 2) and find a roundtrip loss of 1.3%.

Using the above results the cavity decay time is 119ns.

We should investigate what's going on with the ringdown measurements.

Attachment 1: 150MHzSideBandCreation.pdf
150MHzSideBandCreation.pdf
Attachment 2: LinewidthMeasurment.pdf
LinewidthMeasurment.pdf
  15108   Wed Jan 1 04:53:11 2020 gautamUpdatePSLMapping the PSL electronics

For the IMC servo board, it'd be easiest to copy the wiring scheme for the BIO bits as is configured for the CM board (i.e. copy the grouping of the BIO bits on the individual Acromag units). This will enable us to use the latch code with minimal modifications (it was a pain to debug this the first time around). I don't see any major constraint in the wiring assignment that'd make this difficult.

Quote:

PSL wiring spreadsheet is ready. (But the link was stripped. Koji)

Link to a wiki page  with the link to the wiring spreadsheet (Yehonathan)

  15107   Tue Dec 31 03:03:02 2019 gautamUpdatePSLPMC cavity ringdown measurement

When I was looking at this, the AOM shutdown time was measured to be ~120 ns, and while I wasn't able to do a ringdown measurement with the PMC (it'd just stay locked because at the time i was using the zeroth order beam), the PMC transmission decayed in <200 ns. 

  15106   Fri Dec 27 16:26:11 2019 YehonathanUpdatePSLPMC Linewidth measurement

I try to measure the linewidth of the PMC by ramping the PMC PZT. 

I do it by connecting a triangular shape signal to FP Test 1 on the PMC servo front panel (I know, it is probably better to connect it to DC EXT. next time.) and turn the servo gain to a minimum.

Attachment 1 shows the PMC transmission PD as the PZT is swept with the EOM connected and when it is disconnected. It shows the PMC over more than 1 free spectral range.

For some reason, I cannot seem to be able to find the 35MHz sidebands which I want to use to calibrate the PZT scan. I made sure that the EOM is driven by a 35MHz signal using the scope. I also made sure that the PMC cannot to lock without the EOM connected.

I am probably doing something silly.

Attachment 1: PMCTransmissionSpectra.pdf
PMCTransmissionSpectra.pdf
  15105   Fri Dec 27 15:01:02 2019 YehonathanUpdatePSLPMC cavity ringdown measurement

I measured PMC ringdowns for several input powers. I change the input power by changing the DC voltage to the AOM.

First, I raise the DC voltage to the AOM from 0V and observe the signal on the picked off PD. I see that at around 0.6V the signal stops rising. The signal on the PD is around 4V at that point so it is not saturated.

Up until now, we provided 1.5V to the AOM, which means it was saturated.

I measured ringdowns at AOM voltages of 0.05, 0.1, 0.3, 0.5, 1 volt by shutting off the DC voltage to the AOM and measuring the signal at the PMC transmission PD and the picked off PD simultaneously for reference.

Attachment 1 shows the reference measurement for different AOM voltages. For low AOM DC voltages, the response of the AOM+PD is slower.

Attachment 2 shows the PMC transmission PD measurements which barely change as a function of AOM voltage but shows the same trend. I believe that if the AOM+PD response was much faster there would be no observable difference between those measurements.

Attachment 3 shows PMC transmissions and references for AOM voltages 0.05V and 1V. It seems like for low AOM voltages we are barely fast enough to measure the PMC ringdown.

I fitted the 0.3V ringdown and reference to a sum of two exponentials (Attachment 4).

The fitting function is explicitly a * nm.exp(-x/b) +c* nm.exp(-x/d) +e

For the PMC transmission I get:

a = 0.21
b = 3.64 (us)
c = 0.69, 
d = 39.62 (us)
e = 2.0e-04

For the reference measurement:

a = 0.34
b = 4.97 (us)
c = 0.58
d= 31.22 (us)
e = 1.11e-03

I am still not able to do deconvolution of the ref from the measurement reliably. I think we should do a network analyzer measurement.

Shruti, the PD is again in your beam path.

Attachment 1: PDAOMResponse.pdf
PDAOMResponse.pdf
Attachment 2: PMCTransmission.pdf
PMCTransmission.pdf
Attachment 3: RingdownsAndRefs.pdf
RingdownsAndRefs.pdf
Attachment 4: TwoExponentialFitAOM0.3V.pdf
TwoExponentialFitAOM0.3V.pdf
  15104   Mon Dec 23 19:30:20 2019 YehonathanUpdatePSLMapping the PSL electronics

PSL wiring spreadsheet is ready. (But the link was stripped. Koji)

Link to a wiki page  with the link to the wiring spreadsheet (Yehonathan)

  15103   Fri Dec 20 18:33:21 2019 YehonathanUpdatePSLMapping the PSL electronics

Final (hopefully) PSL channel list is attached with allocated Acromag channels. Wiring spreadsheet coming soon.

Current Acromag count:

AI 8
AO 2
BIO 4
Number of channels 8*8+2*8+4*16=144
Number of wires 144*2=288

 

Attachment 1: PSL_Wirings_-_Channel_List.pdf
PSL_Wirings_-_Channel_List.pdf PSL_Wirings_-_Channel_List.pdf PSL_Wirings_-_Channel_List.pdf PSL_Wirings_-_Channel_List.pdf PSL_Wirings_-_Channel_List.pdf
  15102   Tue Dec 17 20:45:30 2019 ranaUpdatePSLPMC cavity ringdown measurement

idk - I'm recently worried about the 'thermal self locking' issue we discussed. I think you should try to measure the linewidth by scanning (with low input power) and also measure the TF directly by modulating the power via the AOM and taking the ratio of input/output with the PDA55s. I'm curious to see if the ringdown is different for low and high powers

Quote:

I plan to model the PD+AOM as a lowpass filter with an RC time constant of 12us and undo its filtering action on the PMC trans ringdown measurement to get the actual ringdown time.

Is this acceptable?

This is an ole SURF report on thermal self-locking that may be of use (I haven't read it or checked it for errors, but Royal was pretty good analytically, so its worth looking at)

  15101   Tue Dec 17 20:08:09 2019 shrutiUpdateGeneralPLL / PM measurement of Xend NPRO PZT

1. Some calculations

For a Unity Gain Frequency (UGF) of 1 kHz, assumed PZT response K_{VCO} of 1 MHz/V, Mixer response K_{M} of 25 mV/\pi rad, the required gain of the amplifier is

G = 2 \pi \times \text{UGF}/ (K_{VCO} K_M)

G ~ 0.8

2. Progress

- Measured the mixer response

Measuring mixer response:

- PSL laser temperature was adjusted so that beat frequency was roughly 25 MHz and the amplitude was found to be roughly -30dBm.

- At the RF port instead of the beat signal, a signal of 25 MHz + few kHz at -30 dBm was inputted. The LO was a 25 MHz signal was sent from the Marconi at 7 dBm.

- The mixer output was measured, with setup as in Attachment 1  Figure (A), on an oscilloscope. The slope near the small angle region of the sine curve would be the gain (in V/rad) and was found to be: K_M \approx 25 \text{ mV}/ \pi rad

- Since from the above calculations it seemed like an amplifer gain of 1 should work for the PLL, I rearranged the set up as in Figure (B) of Attachment 1 to actuate the X end NPRO PZT, I adjusted the PSL temperature (slow control) to try and match the frequency to 25 MHz, but couldn't lock the loop. I was monitoring the error signal after amplification (50 ohm output of the SR 560) which showed oscillations when the beat frequency was near 25 MHz and nothing significant otherwise.

- I used a 20 dB attenuator at the amplifier output and saw the beat note oscillate for longer, but maybe because it was a 50 ohm component in a high impedance channel it did not work either (?). I tried other attenuator combinations with no better luck.

- Is there a better location to add the attenuator? Should I pursue amplifying the beat signal instead?

- Also, it seemed like the beat note drift was higher than earlier. Could it be because the PMC was unlocked?

 

Quote:
 

 

Attachment 1: 20191217.png
20191217.png
  15100   Tue Dec 17 18:05:06 2019 YehonathanUpdatePSLMapping the PSL electronics

Updated the channel list (Attached):

1. Removed the MC steering mirror PZT channels

2. Added Sourcing/Sinking column

3. Recounted the mbbos correctly

4. Allocated Acromags:

Model Purpose No. Spare channels
XT1221 ai 7 11
XT1541 ao + src bo 2 9 ao
XT1121 src bo 2 4
XT1121 sink bo 1 4

I think we can start wiring.

Attachment 1: PSL_Wirings_-_Sheet1_(3).pdf
PSL_Wirings_-_Sheet1_(3).pdf PSL_Wirings_-_Sheet1_(3).pdf
  15099   Tue Dec 17 00:23:28 2019 YehonathanUpdatePSLMapping the PSL electronics

I added to the PSL wiring list the ioo channels and the laser shutter (See attached pdf for an updated list).

The total channel numbers for now:

ai 57
ao 13
bi 1
bo 36

I counted each mbbo as 1 bo but I am not sure that's correct.

Still need to allocate Acromags.

Attachment 1: PSL_Wirings_-_Sheet1_(2).pdf
PSL_Wirings_-_Sheet1_(2).pdf PSL_Wirings_-_Sheet1_(2).pdf
  15098   Mon Dec 16 18:19:42 2019 shrutiUpdatePSLPMC cavity ringdown measurement : beat-note disruption

I have removed the PD55 + ND filter attached to it (see Attachment) and placed it next to the oscilloscope, after disconnecting its output and power supply. The post is still in place.

I did see the beat after that.

Quote:

{Yehonathan, Rana, Jon}

To check whether we laser is being shut fast enough for the ringdown measurement we put a PD55 in the path that leads to the beat note setup. The beam is picked off from the back steering mirror after AOM and before the PMC.

@Shruti the PD is now blocking the beam to your setup.

 

Attachment 1: IMG_0040.jpg
IMG_0040.jpg
  15097   Fri Dec 13 12:28:43 2019 YehonathanUpdatePSLPMC cavity ringdown measurement

I grab the data we recorded yesterday from the scope and plot it in normalized units (remove noise level and divide by maximum). See attachment.

It can be seen that the measured ringdown time is ~ 17us while the shut-off time is ~12us.

I plan to model the PD+AOM as a lowpass filter with an RC time constant of 12us and undo its filtering action on the PMC trans ringdown measurement to get the actual ringdown time.

Is this acceptable?

 

Attachment 1: Ringdown_InitialProcess.pdf
Ringdown_InitialProcess.pdf
  15096   Thu Dec 12 19:20:43 2019 YehonathanUpdatePSLPMC cavity ringdown measurement

{Yehonathan, Rana, Jon}

To check whether we laser is being shut fast enough for the ringdown measurement we put a PD55 in the path that leads to the beat note setup. The beam is picked off from the back steering mirror after AOM and before the PMC.

@Shruti the PD is now blocking the beam to your setup.

As before, we drive the AOM to deflect the beam. The deflected beam is coupled to the PMC cavity. We lock the PMC and then shut the beam by turning off the output of the function generator that provides voltage to the AOM driver.

We measure the transmitted light of the PMC together with the light that is picked off before the PMC. In Attachment 1, the purple trace is the PMC transmission, the green trace is the peaked-off beam and the yellow trace is the function generator signal.

Rana was pointing out that the PDs, the function generator and the scope were not carefully impedance matched, which could lead to erroneous measurements. He also mentioned that the backscattered beam was too bright which might indicate that the PMC is oscillating. To remedy this we lowered the gain of the PMC lock to ~8.

We repeat the measurement after setting all the components to 50ohm (attachment 2). We then realize that the BNC T junction connected on the function generator is splitting the signal between the 50ohm AOM driver and 1Mohm oscilloscope channel which causes distortions as can be seen. We remove the T junction and get a much cleaner measurement (see next elog).

 

It seems like either the shutting speed or the PDs are only slightly faster than the PMC. I also check the AOM driver RF output fall time doing the same kind of measurement (attachment 3).

We suspect the PDs' bandwidth is to blame (although they are quoted to have 10MHz bandwidth).

In any case, this is fast enough for the IMC and arm cavities whose lifetime should be much longer than the PMC.

I will post an elog with some numbers tomorrow.

Attachment 1: IMG_0105.jpeg
IMG_0105.jpeg
Attachment 2: TEK00001.PNG
TEK00001.PNG
Attachment 3: 20191212_151642.jpg
20191212_151642.jpg
  15095   Wed Dec 11 22:01:24 2019 ranaConfigurationComputersMegatron: starts up grade

Megatron is now running Ubuntu 18.04 LTS.

We should probably be able to load all the LSC software on there by adding the appropriate Debian repos.

I have re-enabled the cron jobs in the crontab.

The MC Autolocker and the PSL NPRO Slow/Temperature control are run using 'initctl', so I'll leave that up to Shruti to run/test.

  15094   Wed Dec 11 15:29:17 2019 YehonathanUpdatePSLPMC is locked

Make sure to measure the power drop of the beam downstream of the AOM but before the PMC. Need to plot both together to make sure the chop time is much shorter than the 1/e time.

  15093   Wed Dec 11 15:01:57 2019 JonSummaryPSLPMC cavity ringdown measurement

[Jon, Yehonathan]

We carried out a set of cavity ringdown measurements of the PMC. The 1/e decay time scale is found to be 35.2 +/- 2.4 (systematic) μs. The statistical error is negligible compared to the systematic error, which is taken as the maximum absolute deviation of any measurement from the average value.

To make the measurement, we injected the first order deflection beam of an 80 MHz AOM, then extinguished it quickly by cutting the voltage offset to the AOM driver provided by an RF function generator. A 100 MHz oscilloscope configured to trigger on the falling voltage offset was used to sample the cavity in transmission as sensed by a PDA55. We found the detector noise of the DC-coupled output of the 35.5 MHz REFL PD to be too high for a reflection-side measurement.

Further loss analysis is forthcoming.

Attachment 1: IMG_0101.jpg
IMG_0101.jpg
  15092   Tue Dec 10 18:27:22 2019 YehonathanUpdatePSLPMC is locked

{Yehonathan, Jon}

We are able to lock the PMC on the TEM00 mode of the deflected beam.

However when we turn off the driving voltage to the AOM and back on the lock is not restored. It get stuck on some higher order mode.

There are plethora of modes present when the PZT is scanned, which makes us believe the cavity is misaligned.

 

To lock again on the TEM00 mode again we disconnect the loop (FP Test point 1), find a TEM00 mode using the DC output adjust and close the loop again.

 

  15091   Tue Dec 10 15:17:17 2019 YehonathanUpdatePSLPMC is locked

{Jon, Yehonathan}

We burt-restored the PSL and the PMC locked immediately.

The PMC is now locked on the AOM first order mode.

  15090   Tue Dec 10 13:26:46 2019 YehonathanUpdatePSLAOM first order beam alignment

nvm the PZT can scan over many GHz.

Quote:

 

However, we are currently unable to lock the PMC on high gain. When the gain is too high the PZT voltage goes straight to max and the lock is lost.

Just realized that the diffracted beam is frequency shifted by 80MHz. It would shift the PZT position in the PMC lock acquisition, wouldn't it?

 

  15089   Tue Dec 10 01:24:17 2019 YehonathanUpdatePSLAOM first order beam alignment

 

However, we are currently unable to lock the PMC on high gain. When the gain is too high the PZT voltage goes straight to max and the lock is lost.

Just realized that the diffracted beam is frequency shifted by 80MHz. It would shift the PZT position in the PMC lock acquisition, wouldn't it?

  15088   Mon Dec 9 21:22:46 2019 shrutiUpdateGeneralPLL / PM measurement of Xend NPRO PZT

In short:

Using the same setup as before with a LPF changed to have a cutoff of 5 MHz, the PLL was implemented and a TF measurement of the phase modulation was attempted. But, the beatnote drift was too high to get a prolonged phase lock (many times over 5MHz in <5 min).
 

Steps undertaken:

1. Normally I would unlock the IMC (Disabling the servo between the 'Filter' and 'Polarity' on the Mode Cleaner Servo Screen), but today I did not have to since Rana had kept it unlocked.

2. Misaligned the ITMX. This is to prevent cavity resonances from returning to the laser

3. Turned up the air on the HEPA at the PSL table to 100% during the measurement

4. Cables were connected as before (diagram shown in attachment of elog 15069)

5. The X end laser NPRO was actuated for the TF measurement using a long cable connected to TO AUX_X LASER PZT

 

Thoughts and observations:

- Reading out the error signal after amplification cannot distinguish between a locked loop or one out of its range. The error signal would be very small in both cases.

- Looking at the beat note on an oscilloscope, there also seemed to be an additional amplitude modulation that I had not noticed earlier. Rana suggested that it may have something to do with the pre-mode cleaner and the AOM being driven at 80 MHz

- Even though the TF was attempted, it seemed too noisy, suggesting that the PLL did not seem to work

- Rana also suggested that it may be a better idea to use the PZT of one of the lasers as the VCO for the PLL feedback instead of the Marconi.

 

Quote:

I worked on the setup up for the phase modulation measurement of the X end NPRO PZT. A previous similar measurement can be found here (12077). The setup was assembled based on the schematic in Attachment1.

Mixer used: Level 7, Mini circuits ZP-3+
LPF: up to 1.9MHz


Cables exiting the PSL table:
1. LO (Marconi -> Mixer)
2. RF (PSL+X beat note -> Mixer) The cable for this was taken from the Beat Mouth (otherwise connected to the oscilloscope)
3. Ext modulator (SR560 -> Marconi)

The long cable labled 'X Green Beat' was used to connect to the PZT (from the network analyzer).

Observations: The beat note kept floating between 0 and ~100 MHz

The PLL part of the circuit was tested coarsely with the spectrum analyzer function of the Agilent, where the loop was seen to stabilize when the carrier frequency of the Marconi was close to the instantaneous beat frequency.

 

 

 

  15087   Mon Dec 9 19:19:04 2019 YehonathanUpdatePSLAOM first order beam alignment

{Yehonathan, Rana}

In order to setup a ringdown measurement with perfect extinction we need to align the first order beam from the AOM to the PMC instead of the zeroth order.

We connected a signal generator to the AOM driver and applied some offset voltage. We spot the first order mode and align it to the PMC. The achieved transmitted power is roughly as it was before this procedure.

Along the way few changes has been made in the PSL table:

1. Some dangling BNCs were removed.

2. Laser on the south east side of the PSL table was turned off.

3. DC power supplies were removed (Attachment 1 & 2). The rubber legs on the first one are sticky and leave black residue.

4. The beam block that orginally blocked the AOM high order modes was raised to block the zeroth order mode (Attachment 3).

5. The unterminated BNC T junction (Attachment 4 - before picture). from the PMC mixer to the PMC servo was removed.

However, we are currently unable to lock the PMC on high gain. When the gain is too high the PZT voltage goes straight to max and the lock is lost.

Attachment 1: 20191209_193112.jpg
20191209_193112.jpg
Attachment 2: 20191209_193203_HDR.jpg
20191209_193203_HDR.jpg
Attachment 3: imageedit_2_7551928142.gif
imageedit_2_7551928142.gif
Attachment 4: imageedit_3_5863650538.gif
imageedit_3_5863650538.gif
  15086   Mon Dec 9 13:08:24 2019 YehonathanSummaryPEMJump test of seismometers: EX needs recentering

I check the seismometers in the last 14 hours (Attached). Seems like the coherenece is restored in the x direction.

 

 

Attachment 1: seis_191208.pdf
seis_191208.pdf
  15085   Sun Dec 8 20:48:29 2019 ranaConfigurationComputersMegatron: starts up grade

I noticed recently that Megatron was running Ubuntu 12, so I've started its OS upgrade.

  1. Unlocked the IMC + disabled the autolocker from the LockMC screen + closed the PSL shutter (IMC REFL shutter doesn't seem to do anythin)
  2. Disabled the "FSS" slow servo on the FSS screen
  3. did sudo apt-get update, sudo apt-get upgrade, and then sudo apt-get do-release-upgrade which starts the actual thing
  4. According to the internet, the LTS upgrades will go in series rather than up to 18 in one shot, so its now doing 12 -> 14 (Trusty Tapir)

Megatron and IMC autolocking will be down for awhile, so we should use a different 'script' computer this week.


Mon Dec 9 14:52:58 2019

upgrade to Ubuntu 14 complete; now upgrading to 16

  15084   Sun Dec 8 20:27:11 2019 ranaUpdateComputersViviana upgrade to Ubuntu 16

The IBM laptop at EX was running Ubuntu 14, so I allowed it to start upgrading itself to Ubuntu16 as it desired. After it is done, I will upgrade it to 18.04 LTS. We should have them all run LTS.

  15083   Sun Dec 8 20:15:41 2019 ranaSummaryPEMJump test of seismometers: EX needs recentering

I have re-centered the EX (and EY) seismometers. They are Guralp CMG40-T, and have no special centering procedure except cycling the power a few times. I turned off the power on their interface box, then waited 10s before turning it back on.

The fist atm shows the comparison using data from 8-9 PM Saturday night:

  1. there seems to be a factor of 2 calibration diff between the T240 near the BS, and the Guralp seismometers at the end. Which one is right? surpriseWhen was the last time they were cross calibrated?
  2. The low coherence between BS_X and EX_X shows the problem. They should be very coherent (> 0.9) for 0.1-1 Hz.sad

 

Attachment 1: seis_all_191208.pdf
seis_all_191208.pdf
  15082   Fri Dec 6 17:49:46 2019 ranaSummaryPEMJump test of seismometers: EX needs recentering

Yehonathan, please center the EX seismometer.

The attached PDF shows the seismometer signals (I'm assuming that they're already calibrated into microns/s) during the lab tour for the art students on 11/1. The big spike which I've zoomed in on shows the time when we were in the control room and we all jumped up at the same time. There were approximately 15 students each with a mass of ~50-70 kg. I estimate that out landing times were all sync'd to within ~0.1 s.

Attachment 1: Seismometers.pdf
Seismometers.pdf
Attachment 2: src.tgz
  15081   Fri Dec 6 15:22:01 2019 gautamFrogsLSCDAFI system revived

[Jordan, gautam]

We did the following:

  • Route the fiber from the control room to 1Y2.
  • Plug fiber in to FiBox at either end, turned FiBoxes ON.
  • Tested the optical connection by driving a 1Vpp 440 Hz sine wave from a function generator - Yehonathan hears it loud and clear in the control room.
  • Tested that both CH1 and CH2 work - only CH1 is connected to the speakers in the control room at the moment.
  • There is some cross-coupling between the channels - not sure if this is happening in the multi-mode fiber or in the electroncis, but I estimate the isolation to be >30dB.
  • Connected CH8 and CH9 of DAC0 in the c1lsc expansion chassis to CH1 and CH2 respectively of the FiBox in 1Y2. 
  • Restarted the c1daf model on c1lsc, came up smooth.
  • Routed the POY11 error signal through the various matrices in c1daf, and we could 👂 the Y-arm cavity 🔐 😎 
  • Channels are muted for now - I'll give this a whirl while doing the PRFPMI locking.
  15080   Fri Dec 6 00:02:48 2019 gautamUpdateLSCWhat is the correct way to set the 3f offsets?

Summary:

I made it to 0 CARM offset, PRMI locked a bunch of times today. However, I could not successfully engage the AO path.

Details:

Much of the procedure is scripted, here is the rough set of steps:

  • Transition control of the arms from IR signals to ALS signals.
  • DC couple the ITM oplev servos
  • Burt-restore the settings for PRMI locking with REFL165I-->PRCL, REFL165Q-->MICH, and then enable the MICH_B / PRCL_B locking servos.
  • Add some POPDC to the PRMI triggering (nominally only POP22_I) to let these loops be locked while POP22_I fluctuates wildly when we are near the CARM=0 point.
  • Zero the CARM offset.
  • Adjust the CARM_A/DARM_A offsets such that CARM_B/DARM_B are fluctuating symmetrically about 0.
  • CARM_B gain --> 1.0, to begin the RF blend.
  • Prepare to hand the DC control authority to ALS by turning off FM1 in the CARM filter bank, and turning ON an integrator in the CARM_B filter.

As I type this out, I realized that I was incorrectly setting offsets to maximize the arm powers by adjusting CARM/DARM offsets as opposed to CARM_A / DARM_A offsets. Tried another round of locking, but this time, I can't even turn the integrator on to get the arms to click into somewhat stable powers. 

One thing I noticed is that depending on the offsets I put into the 3f locking loops, the mean value of REFL11 and AS55 when the ALS CARM/DARM offsets are zeroed changes quite significantly. What is the correct condition to set these offsets? They are different when locking the PRC without arm cavities, and also seem to change continuously with CARM offset. I am wondering if I have too much offset in one of the vertex locking loops?

  15079   Thu Dec 5 18:15:01 2019 gautamUpdateOptical LeversITM, PRM and BS Oplevs re-centered

In preparation for locking tonight, I re-centered the spots on the Oplev QPDs for the ITMs, BS and PRM after locking and running the dither alignment for the arms and also the PRMI carrier. In the past, DC coupling the ITM Oplevs helped the angular stability a bit, let's see if it still does.

  15078   Thu Dec 5 15:09:50 2019 gautamUpdateCDSc1oaf crashed c1lsc

I tried starting the c1oaf model, but got a DQ error (I want the option of running feedforward during locking even if the filters aren't particularly well tuned yet). Note that this isn't "just a warning light" - some channels are initialized to +/- 1e20, so if you try turning some filters on, you will deliver a massive kick to the optics. Restarting it crashed c1lsc (this is not unexpected behavior - the only way to clear the DQ error is to restart the model, and empirically, the success rate is ~50%). The reboot script brought everything back online smoothly, and the second, time, c1oaf started without any issues.

While looking at the CDS overview screen, I noticed that the c1scy model was reporting frequent RFM errors for the C1:SCY-RFM_ETMY_LSC channel (but none of the others). On the sender model (c1rfm), no errors were being reported. The diag reset button / mxstream restart didn't really work either. See Attachment #1. Just restarting the c1scy model didn't fix the error - I had to reboot the machine and restart the models, and now no errors are being reported.

Attachment #2 shows the current nominal CDS status - the red light on c1lsc is due to some missing c1dnn channels (I'll remove these at the next c1lsc model change because I don't want to un-necessarily reboot the vertex FEs), and the c1omc model is obsolete I guess. c1daf isn't running right now but once I get the new fiber (ordered), I'm gonna restart this model as well.

P.S. The ALS temperature sliders are not SDF-ed. So when the model was restarted, I had to change the sliders back to their old values to get the beat back in the usable range.

Attachment 1: SCYerrors.png
SCYerrors.png
Attachment 2: CDSnormal.png
CDSnormal.png
  15077   Thu Dec 5 14:54:15 2019 gautamUpdateGeneralSymlink to SRmeasure and AGmeasure

I symlinked the SRmeasure and AGmeasure commands to /usr/bin/ on donatella (as it is done on pianosa) so that these scripts are in $PATH and may be run without having to navigate to the labutils directory.

  15076   Thu Dec 5 08:44:44 2019 GavinUpdateLoss MeasurementQ Measurements of Test Masses

[Yehonathan, Gavin]

Measuring POX11_Q_MON and injecting a signal into the ITMX_UL_IN port a signal could not be seen on the function generator. After debugging the source of the issue was two fold:

  • By using the quadrant drives for coils (UL, UR etc) a signal has to pass through a switch before reaching the driver. To resolve this the signal input was switched to POS_IN (driving the entire coil at once rather than in quadrants) which has no switch to bypass.
  • The averaging on the Stanford SR785 was set too low. By increasing the averages from 10 to 25 the signal became more visible.

Unrelated to these issues the signal input was switched to POY11_Q_MON and ITMY_POS_IN as part of the debugging process. The function generator used was switched from the Stanford to the Siglant SDG 1032X.

An unrelated issue but note worthy was the Lenovo 40m laptop used to measure the IFO state (locked or unlocked) ran out of battery in a very short timespan.

To gauge where the resonance of the test masses are FEA model of a simple 40m test mass was computed to give an esitimate at what frequency the eigenmodes exist. For the first two modes the model gave resonances at 20.366 kHz (butterfly mode) and 28.820 kHz (drumhead mode). Then by measuring with an acquisition time of 1 s at they frequencies on the SR785 and injecting broad band white noise with a mean of 0 V and a stdev of 2 V, small peaks were seen above the noise at 20.260 kHz and 28.846 kHz. By then injecting a sine wave at those frequencies with 9 Vpp, the peak became clearly visible above the noise floor.

The last step is to measure the natural decay of these modes when the excitation is turned off. It is difficult to tell currently if these are indeed eigenmodes or just large cavity injections with an associated stabilisation time (what could appear as a ringdown decay). More investigation is required.

 

Attachment 1: 20191205_132158.jpg
20191205_132158.jpg
  15075   Thu Dec 5 01:54:39 2019 gautamUpdateLSCPartial CM board path engaged
  • The arm powers could be stabilized somewhat once the CM_SLOW path to MC2 was engaged.
  • However, I was never able to get the AO path to do anything good.
  • Took a bunch of CM board TFs, need to think about what I need to do differently to get this next bit to work.
  • An SR785 is sitting next to the LSC rack hooked up to the CM board. I also borrowed the GPIB unit from the AG4395 to grab data from said SR785.
  • One thing I noticed that the CARM_B (=CM_SLOW) and DARM_B (=AS55_Q) signals both had a DC offset, so maybe this is indicative of some DC offset in the PRMI 3f signals? Right now, I lock the PRMI without any offsets, and as I reduce the CARM offset, I can see the DC value of REFL11_I and AS55_Q changing significantly. To be investigated in tonight's locking.
Attachment 1: AOengaged.pdf
AOengaged.pdf
  15074   Wed Dec 4 20:32:43 2019 gautamUpdateGeneralPLL for PM measurement

Were some cables from the ALS beat setup modified? I can't see the beat on the scope, and this elog doesn't say anything about cable connection rearrangement. At ~2311, I am reverting the setup to as it should be.

  15073   Wed Dec 4 19:54:27 2019 gautamUpdateLSCA look to the past

Trawling through some past elogs, I saw that the ALS noise increase as a function of CARM offset reduction is not really a new thing (see e.g. this elog). In the past, when we were able to lock, when the CARM offset is reduced to zero, the arms would "buzz" through resonance. It just wasn't clear to me how much the buzzing was - in all the plots we presented, we were not looking at the fast 16k output, so it looked like the arm powers had stabilized. But today, looking at the frame data at 16k from back in 2016, it is clear to me that the arm transmission was in fact swinging all the way from 0 to some maximum. Once the IR signal (=REFL11) blending is turned on, we were able to stabilize the arm power somewhat. What this means is that we are in a comparable state as to when we were able to lock in the past (since I'm able to sit at 0 CARM offset with the PRMI locked almost indefinitely).

So, I think what I'll try for the next 3 days is to get this blending going, I think I couldn't enable the CM_slow path because when I was experimenting with the high bandwidth Y arm cavity locking, I had increased the whitening gain of this channel, but REFL11 has much more optical gain (=larger signal) than POY11, and so I'll start from 0dB whitening gain and see if I can turn the magic integrator on. Long term, we should try and compensate the optomechanical plant that changes as our CARM offset gets reduced, as this would further reduce the lock acquisition time and simplify the procedure (no need to fiddle with the integrator, offsets etc). A relevant thread from the past.

Attachment 1: DRFPMI_2016March.pdf
DRFPMI_2016March.pdf
  15072   Wed Dec 4 12:13:10 2019 gautamUpdateCDSReboot script

It was way more annoying without a script and took longer than the 4 minutes it does now.

You can fix the requirement to enter password by changing the sshd settings on the FEs like I did for pianosa.

After running the script, you should verify that there are no red flags in the output to console. Yesterday, some of the settings the script was supposed to reset weren't correctly reset, possibly due to python/EPICS problems on donatella, and this cost me an hour of searching last night because the locking wasn't working. Anyway, best practise is to not crash the FEs.

Quote:

The script is a bit annoying in that it requires entering the CDSs' passwords multiple times over the time it runs which is long.

  15071   Wed Dec 4 09:11:42 2019 YehonathanUpdateCDSReboot script

After the CDSs crashed we run the rebootC1LSC.sh script.

The script is a bit annoying in that it requires entering the CDSs' passwords multiple times over the time it runs which is long.

The resulting CDS screen is a bit different than what was reported before (attached). Also, not all watchdogs were restored.

We restore the remaining watchdogs and do XARM locking. Everything seems to be fine.

Attachment 1: medmScreen11.ps
medmScreen11.ps
  15070   Wed Dec 4 08:54:07 2019 YehonathanUpdateIMCMirror analog shaking

{Yehonathan, Gavin}

Yesterday we tried to shake ITMX with a function generator in order to observe the 28.8kHz drum mode.

We laid a long BNC cable that runs from the YARM to the XARM. This cable either needs to be collected back to the BNC big plastic cable box under the IMC or be labeled so that it could be found easily in the future.

First, we tried to shake it at a lower frequency (100's of Hz) where the shaking should be easily observed in the POSX channel. We try driving the POS channel on the ITMX servo but nothing happens. Most likely it is disconnected.

While setting up for shaking the individual OSEM channels 4 CDSs crashed (c1lsc, c1ass, c1oaf, c1cal).

 

  15069   Tue Dec 3 22:41:17 2019 shrutiUpdateGeneralPLL for PM measurement

I worked on the setup up for the phase modulation measurement of the X end NPRO PZT. A previous similar measurement can be found here (12077). The setup was assembled based on the schematic in Attachment1.

Mixer used: Level 7, Mini circuits ZP-3+
LPF: up to 1.9MHz


Cables exiting the PSL table:
1. LO (Marconi -> Mixer)
2. RF (PSL+X beat note -> Mixer) The cable for this was taken from the Beat Mouth (otherwise connected to the oscilloscope)
3. Ext modulator (SR560 -> Marconi)

The long cable labled 'X Green Beat' was used to connect to the PZT (from the network analyzer).

Observations: The beat note kept floating between 0 and ~100 MHz

The PLL part of the circuit was tested coarsely with the spectrum analyzer function of the Agilent, where the loop was seen to stabilize when the carrier frequency of the Marconi was close to the instantaneous beat frequency.

 

 

Attachment 1: PM_measurement.jpeg
PM_measurement.jpeg
ELOG V3.1.3-